Find all values of x in the interval [0, 2π] that satisfy the given equation. (Enter your answers as a comma-separated list.)
(sin x)2 − 6 cos x − 6 = 0

Answer by lwsshak3(11628)

Find all values of x in the interval [0, 2π] that satisfy the given equation. (Enter your answers as a comma-separated list.)
(sin x)2 − 6 cos x − 6 = 0
**
(sin^2x − 6 cos x − 6 = 0
1-cos^2x-6cosx-6=0
cos^2x+6cosx+5=0
(cosx+1)(cosx+5)=0
cosx+5=0
cosx=-5 (reject, -1 ≤ cosx ≤ 1
cosx+1=0
cosx=-1
x=π